OpenStudy (anonymous):

a ballon with a basket with a man inside starts from rest and rises vertically with a uniform acceleration of 4.9m/s2 the man releases a ball 2 second after the ballon leaves the ground the maximum height above the ground attained by the ball is???

OpenStudy (experimentx):

calculate final velocity, v = u + at , u=0, t=2, a=4.9 and use the prevous v as u here, v^2 = u^2 + 2as, v=0, a = -g, find s

OpenStudy (anonymous):

plz can u solve it

OpenStudy (anonymous):

option are (a) 9.8 (b) 14.7 (c) 19.6 (d) 24.5 and the answer is b can u solve it

OpenStudy (experimentx):

v= 9.8 => u = 9.8 0 = 9.8^2 - 2* 9.8 * s => s = 9.8/2 = 4.9 m Looks like we also have to consider the distance traveled at 2s before s= ut + 1/2 at^2 = 0 + 1/2 * 4.9*2^2 = 9.8 m add those you get, 4.9 + 9.8 and get your answer.

OpenStudy (anonymous):

s=ut+1/2at\[t ^\] s=0x2+1/2x10x4 s=0+20 s=20m